OpenStudy (anonymous):

Cot theta = 0.4986 whats angle theta?

OpenStudy (shane_b):

\[cos^{-1}(0.4986)=?\]

OpenStudy (anonymous):

my answer choices are a. 26 degree 30' b. 60 degree 10' c. 63 degree 30' d. 64 degree 30' e. 29 degree 50'

OpenStudy (anonymous):

what about sin theta= 0.0987

OpenStudy (matheducatormcg):

i think the original question is about cot not cos.

OpenStudy (shane_b):

Ack...that makes sense why my answer was slightly off :P

OpenStudy (anonymous):

so is it still b?

OpenStudy (shane_b):

\[cot \theta=\frac{1}{tan{\theta}}=0.4986\]\[\frac{1}{0.4986}=tan \theta\]\[tan^{-1}(\frac{1}{0.4986})=63.499\] That's about 64 degrees and 30'. Sorry...had a phone call.

OpenStudy (shane_b):

skfaljkhfas 63 degrees, 30 minutes. I must be tired. Answer C!

OpenStudy (anonymous):

Ok, thanks. And how about sin theta= 0.0987?

OpenStudy (matheducatormcg):

apply inverse trigonometry. trig functions typically get radians or degrees as inputs. using inverse trig functions, the value that the trig function is equal to becomes your input. i even had inverse trig functions on my last cell phone. google chrome will usually solve problems if you enter them on the address bar.

OpenStudy (matheducatormcg):

just typed in arcsin(.0987)=degrees and it gave me 5.664315618

OpenStudy (matheducatormcg):

typed into google chrome address bar that is.